# Six-Point Method for Multi-Camera Systems with Reduced Solution Space

### v1

@@ 5 Conclusion
We proposed a series of minimal solvers to compute the 6DOF relative pose of a multi-camera system using a minimal number of six PCs. We also exploit ray bundle constraints that allow for a reduction of the solution space and the development of more stable solvers. A generic solver is proposed for the relative pose estimation of generalized cameras. All configurations of minimal six-point problems for multi-camera systems are enumerated. Moreover, two minimal solvers, including an inter-camera solver and an intra-camera solver, are proposed for practical configurations of two-camera rigs. Based on synthetic and real-world experiments, we demonstrate that the proposed solvers offer an efficient solution for ego-motion estimation, surpassing state-of-the-art solvers in terms of accuracy.
Supplementary Material
6 Proof of a Common Factor
In Section 3.3 of the paper, we say that a factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ can be factored out in Eqs. (16) and (17) of the paper. This property can be proven by symbolic computation using Matlab. The Matlab code is listed in LABEL:lst:proof.
@@ 6 Proof of a Common Factor
It can be seen that $3\times 3$ and $4\times 4$ submatrices of ${\mathbf{M}}(q_{x},q_{y},q_{z})$ in Eq. (9) of the paper have a common factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$. Thus, we can use this property to factor out $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ to simplify the equation system in the paper. It will generate more efficient solvers and sometimes avoid false roots.
7 Configuration Enumeration for Six-Point Solvers
In Section 3.4 of the paper, two configurations called inter-camera and intra-camera are investigated for the six-point problem for two-camera rigs. Please refer to Fig. 3 in the paper. Then a question naturally appears as follows :
@@ 7 Configuration Enumeration for Six-Point Solvers
7.1 Problem Definition
A configuration of point correspondences (PCs) can be modeled by a directed graph (also called digraph) by the following procedure. The vertex $i$ in the graph represents the single camera $i$ in a multi-camera system. An edge from vertex $i$ to $j$ in the graph represents a PC from camera $i$ in the first view to camera $j$ in the second view. For example, the inter-camera and intra-camera cases in the Fig. 3 in the paper can be represented by Fig. 6. Then enumerating all the configurations can be equivalently converted to a graphical enumeration problem [18].
@@ 7.1 Problem Definition
Problem Reformulation: We aim to count the $n$-edge directed graphs. The graphs do not have an isolated vertex. The graphs can contain loops (edges connecting a vertex to itself) and multiple edges (more than one edge connecting the same pair of vertices). How to count and enumerate distinct graphs for a given edge number $n$? Two graphs are distinct if and only if two conditions are met: (1) they are not isomorphic, (2) they cannot be converted to each other by reversing all edge directions of one graph. A graph $\mathcal{G}$ consists of a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of vertices and $\mathcal{E}$ the set of edges. In the paper, we are particularly interested in $n=6$, i.e., a minimal configuration of six PCs.
Isomorphism [Section 5 of [17]]: Two graphs $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ and $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ are isomorphic if there is a bijection $f:\mathcal{V}_{1}\to\mathcal{V}_{2}$ such that a directed edge $(u,v)\in\mathcal{E}_{1}$ if and only if $(f(u),f(v))\in\mathcal{E}_{2}$. In addition, the repetition numbers of $(u,v)$ and $(f(u),f(v))$ are the same if multiple edges or loops are allowed. This bijection $f$ is called an isomorphism.
@@
In the reformulated problem, the graph isomorphism is used to exclude the symmetry of camera indices. In other words, swapping camera indices do not change the configuration. The edge-reversion symmetry is used to exclude the symmetry of two views. A directed graph $\mathcal{G}$ with edges =$\mathcal{E}=\{(u_{i},v_{i})\}_{i=1}^{n}$ and no isolated vertex can be represented by
@@
It is obvious that rearranging columns does not change the graph at all. In addition, a permutation of vertices in $E$ causes an isomorphic graph. Swapping two rows corresponds to reversing edge directions of this graph. In this paper, we write $E_{1}\cong E_{2}$ if $E_{2}$ can be obtained by rearrangement of columns, swapping rows, and swapping vertices of $E_{1}$.
@@
It can be verified that $E_{3}$ can be convert to $E_{2}$ by swapping vertices $1$ and $3$; $E_{6}$ can be convert to $E_{2}$ by swapping vertices $1$ and $2$. $E_{5}$ can be convert to $E_{4}$ by swapping its two rows. As a result, there are $3$ distinct graph. The distinct graphs with isomorphism and edge-reversion symmetry are $E_{1}$, $E_{2}\cong E_{3}\cong E_{6}$, and $E_{4}\cong E_{5}$.
7.2 A Combinatorics Solution to Enumeration
Table 5: The number $D_{n}$ of distinct graphs given $n$ edges.
#edge $n$ 1 2 3 4 5 6 7 8 9 10
#distinct graph 2 9 37 186 985 5953 38689 271492 2016845 15767277
#camera 1 2 3 4 5 6 7 8 9 10 11 12
#cases 1 29 270 1029 1776 1630 853 280 66 15 3 1
#camera 1 2 3 4 5 6 7 8 9 10 11 12 13 14
#cases 1 39 568 3316 8599 11516 8787 4170 1296 312 66 15 3 1
The problem is related to combinatorics and group theory. When isomorphism without edge-reversion symmetry is considered in multigraph enumeration, the problem has been investigated in [41]. However, when edge-reversion symmetry is also considered, little study has been investigated. The Pólya Enumeration Theorem can be applied to solve this problem [17]. To apply this theorem, we need to generate the cycle index of the permutation group $S_{n}$ acting on $n$ elements. Then we obtain the number of isomorphic graphs from this cycle index of $S_{n}$.
@@ 7.2 A Combinatorics Solution to Enumeration
Listing 2: Generate Distinct Graphs Using Matlab
@@
7.3 Statistics of Cases and Solvers
  • Match type $6\cup\phi$: All the 6 edges are same. There are $2$ cases belong to this match type, including
    $E_{1}=\left[\begin{array}[]{cccccc}1&~1&~1&1&1&1\\ 1&~1&~1&1&1&1\end{array}\right],$
    $E_{2}=\left[\begin{array}[]{cccccc}1&~1&~1&1&1&1\\ 2&~2&~2&2&2&2\end{array}\right].$
    These two cases essentially correspond to monocular cameras. Since $6$ PCs provide an over-determined equation system, the equation system has none solution.
    $E_{3}=\left[\begin{array}[]{cccccc}1&~1&~1&1&1&2\\ 1&~1&~1&1&1&3\end{array}\right].$
    $E_{4}=\left[\begin{array}[]{cccccc}1&~1&~1&1&2&4\\ 1&~1&~1&1&3&5\end{array}\right].$
  • Match type $3\cup 3$: The maximal repetitive number of multiple edges is $3$, and there are two multiple edges satisfying this condition. For example, the previously defined inter-camera and intra-camera cases belong to this match type. They are
    $E_{5}=\left[\begin{array}[]{cccccc}1&~1&~1&2&2&2\\ 2&~2&~2&1&1&1\end{array}\right],$
    $E_{6}=\left[\begin{array}[]{cccccc}1&~1&~1&2&2&2\\ 1&~1&~1&2&2&2\end{array}\right].$
  • Match type $3\cup x$: The maximal repetitive number of multiple edges is $3$, and there are only one multiple edge satisfying this condition. For example,
    $E_{7}=\left[\begin{array}[]{cccccc}1&~1&~1&2&4&6\\ 1&~1&~1&3&5&7\end{array}\right].$
  • Match type $\cup_{i}x_{i}$: The maximal repetitive number of multiple edges is less or equal than $2$. For example,
    $E_{8}=\left[\begin{array}[]{cccccc}1&~3&~5&7&9&11\\ 2&~4&~6&8&10&12\end{array}\right].$
For different match types, the numbers of equations in $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are different. We found that cases belong to the same match type have preciously the same number of solutions. Tab. 8 shows the number of solutions for different match types in six-point problems for multi-camera systems.
@@ 7.3 Statistics of Cases and Solvers
8 Experiments
@@
In this section, we add more experiments about the 6DOF relative pose estimation of a multi-camera system. As described in the paper, all the solvers are implemented in C++ [^2] Source codes are provided in the supplementary material and will be released.. The source codes of the comparison methods are taken from the OpenGV library [23]. The numerical stability of the proposed solvers under the second simulated scenario is shown in Sec. 8.1. Expected runtime vs. inlier ratio accounting for numerical stability is demonstrated in Sec. 8.2. More experiments on synthetic data are shown in Sec. 8.3. The real-world experiments on KITTI, nuScenes, and EuRoC datasets are shown in Sec. 8.4, Sec. 8.5 and Sec. 8.6, respectively.
8.1 Numerical Stability
@@
In Section 4.1 of the paper, we design two simulated scenarios to test the numerical stability. First, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras. Second, we design a simulated generalized camera composed of omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. We report the results for the solvers under the first scenario below in the paper. The results for generalized cameras under the second scenario are shown in the supplementary material.
@@
Fig. 7 reports the numerical stability of the solvers on noise-free observations under the second scenario. The numerical stability of the solvers is tested in generic-camera, inter-camera, and intra-camera configurations. The strategy of correspondence selection is described in Sec. 8.3.2. For all three cases, the proposed solvers have satisfactory numerical stability. Among them, the 6pt-Our-inter56 solver has the best numerical stability. None of the comparison solvers can have satisfactory results for all the cases. Specifically,
@@
  • 6pt-Stewénius has satisfactory results for generic-camera configuration. However, due to their bad numerical stability, they do not work well for inter-camera and intra-camera cases. This phenomenon is consistent with [22], which observes the 6pt-Stewénius solver does not work for most axial cameras.
(a) Rotation error $\varepsilon_{\mathbf{R}}$.
(b) Translation error $\varepsilon_{\mathbf{t}}$.
(c) Translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$.
(d) Rotation error $\varepsilon_{\mathbf{R}}$.
(e) Translation error $\varepsilon_{\mathbf{t}}$.
(f) Translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$.
(g) Rotation error $\varepsilon_{\mathbf{R}}$.
(h) Translation error $\varepsilon_{\mathbf{t}}$.
(i) Translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$.
(j) Rotation error $\varepsilon_{\mathbf{R}}$.
(k) Translation error $\varepsilon_{\mathbf{t}}$.
(l) Translation direction error $\varepsilon_{\mathbf{t},\text{dir}}$.
Figure 7: Probability density functions over pose estimation errors on noise-free observations for the proposed solvers. We design a simulated multi-camera system composed of twelve omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. The horizontal axis represents the $\log_{10}$ errors, and the vertical axis represents the density.
@@
Figure 8: RANSAC iteration number with respect to outlier ratio for success probability $99\%$.

### v2

@@ 5 Conclusion
We proposed a series of minimal solvers to compute the 6DOF relative pose of multi-camera systems using a minimal number of six PCs. We also exploit ray bundle constraints that allow for a reduction of the solution space and the development of more stable solvers. A generic solver is proposed for the relative pose estimation of generalized cameras. All configurations of minimal six-point problems for multi-camera systems are enumerated. Moreover, two minimal solvers, including an inter-camera solver and an intra-camera solver, are proposed for practical configurations of two-camera rigs. Based on both synthetic and real-world experiments, we demonstrate that the proposed solvers offer an efficient solution for estimating ego-motion of multi-camera systems, surpassing state-of-the-art solvers in terms of accuracy.
Acknowledgments
This research has been supported in part by the Hunan Provincial Natural Science Foundation for Excellent Young Scholars under Grant 2023JJ20045, and the National Natural Science Foundation of China under Grant 12372189. Further funding support is provided by project 62250610225 by the Natural Science Foundation of China, as well as projects 22DZ1201900, 22ZR1441300, and dfycbj-1 by the Natural Science Foundation of Shanghai. J. Zhao would like to thank Saibal Mitra, from the Netherlands, for the valuable discussions on graph enumeration.
Supplementary Material for
Six-Point Method for Multi-Camera Systems with Reduced Solution Space
[…]
Appendix
Appendix 0.A Related Work
Given that each affine correspondence yields three geometric constraints, a minimal solver using two affine correspondences was proposed to estimate the 6DOF relative pose [14], which is related to our method in this paper. Although both methods use the classical hidden variable technique to decouple rotation and the other variables, and use Gröbner basis technique to find algebraic solutions, we clarify that there are three significant differences as follows: (1) Different parametrizations of the relative pose problem are used, which leads to different polynomial equation systems. A further fundamental difference lies in the geometric constraints derived from point or affine correspondences. Even if the highest-degree monomial is the same, the polynomial equation systems are still different. (2) Finding practical solvers for the six-point problem using the rotation-translation parametrization is certainly non-trivial and there are no off-the-shelf solvers. This requires instantiating random problems in a finite prime field while preserving all geometric side constraints. (3) Two contributions are newly proposed. The ray bundle constraints are proven and exploited for solution space reduction. The ray bundle shares a common origin and propagates in different directions, which is different from the implicit geometric constraints in [14]. Moreover, all configurations of minimal six-point problems for multi-camera systems are first enumerated using the Pólya enumeration theorem.
Appendix 0.B Proof of a Common Factor
In Section 3.3 of the paper, we say that a factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ can be factored out in Eqs. (14) and (15) of the paper. The existence of a common factor can be proven through formula derivation [53]. In this paper, we show that this property can be also proven by symbolic computation using Matlab. The Matlab code is listed in LABEL:lst:proof.
@@ Appendix 0.B Proof of a Common Factor
It can be seen that $3\times 3$ and $4\times 4$ submatrices of ${\mathbf{M}}(q_{x},q_{y},q_{z})$ in Eq. (8) of the paper have a common factor $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$. Note that representing all variables with symbols, that is, changing random numbers in the code to symbols, makes a rigorous proof. However, the runtime of symbolic computation will become about $15$ hours.
Based on the existence of the common factor, we can use this property to factor out $q_{x}^{2}+q_{y}^{2}+q_{z}^{2}+1$ which simplifies the equation system. It leads to the generation of more efficient solvers while also potentially avoiding false roots.
Appendix 0.C Six-Point Solvers for Generalized Cameras
In Section 3.4 of the paper, we discuss the application of the six-point method to the generalized camera model, which is a concept larger than multiple camera systems. Specifically, we show that six-point matches of a generalized camera can be equivalently captured by virtual 12-perspective-camera rigs, making the six-point method applicable to any generalized camera.
(a) $i$-th match of a generalized camera.
(b) virtual 2-perspective-camera rig for $i$-th match.
Figure 5: Application of the six-point method to the generalized camera model.
A point match of a spatial point captured by a generalized camera can be described by two rays. As shown in Fig. 5(a), the intersection of two rays demonstrates that the same spatial point has been observed by a generalized camera in two views. For the $i$-th point match, denote $o_{i}$ and $o^{\prime}_{i}$ as points lying on these two rays. Then, we can define a virtual 2-perspective-camera rig whose reference is consistent with the original generalized camera, as shown in Fig. 5(b). The optical centers of the perspective cameras are $o_{i}$ and $o^{\prime}_{i}$, expressed in the generalized camera reference, respectively. The orientations of the virtual cameras are somewhat arbitrary since the virtual cameras can have a large FOV to capture the spatial point for the $i$-th match. Note that the extrinsic parameters of the virtual cameras are not coupled with the unknown relative pose of the generalized camera.
Based on the previous analysis, we have the following conclusions: (1) Each point match of a generalized camera can be equivalently captured by two virtual perspective cameras. (2) The construction (and thus the extrinsic parameters) of the virtual cameras is not unique. For possible virtual cameras, their observations for the point match are determined by corresponding extrinsic parameters. As a result, the extrinsic parameters of virtual cameras will not change the relative pose estimation results. (3) Since we focus on six-point matches, a virtual 12-perspective-camera rig should be constructed. Then the proposed generic six-point solver can be applied.
Appendix 0.D Configuration Enumeration for Six-Point Solvers
In Section 3.4 of the paper, two practical configurations called inter-camera and intra-camera are investigated for the six-point problem for two-camera rigs. Please refer to Fig. 3 in the paper. Then a question naturally appears as follows:
@@ Appendix 0.D Configuration Enumeration for Six-Point Solvers
0.D.1 Problem Definition
A configuration of point correspondences (PCs) can be modeled by a directed graph (also called a digraph) by the following procedure. The vertex $i$ in the graph represents the single camera $i$ in a multi-camera system. An edge from vertex $i$ to $j$ in the graph represents a PC from camera $i$ in view 1 to camera $j$ in view 2. For example, the inter-camera and intra-camera cases in Fig. 3 in the paper can be represented by Fig. 6. Then enumerating all the configurations can be equivalently converted to a graphical enumeration problem [18].
@@ 0.D.1 Problem Definition
Problem Reformulation: We aim to count the $n$-edge directed graphs. The graphs do not have an isolated vertex. The graphs allow for the presence of loops, which denote edges connecting a vertex to itself, and multiple edges, indicating the existence of more than one edge connecting the same pair of vertices.
How to count and enumerate distinct graphs for a given edge number $n$? Two graphs are distinct if and only if two conditions are met: (1) they are not isomorphic, (2) they cannot be converted to each other by reversing all edge directions of one graph. A graph $\mathcal{G}$ comprises a pair $(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents the set of vertices, and $\mathcal{E}$ represents the set of edges. In this paper, we are particularly interested in $n=6$, i.e., a minimal configuration of six PCs.
Isomorphism [Section 5 of [17]]: Two graphs $\mathcal{G}_{1}=(\mathcal{V}_{1},\mathcal{E}_{1})$ and $\mathcal{G}_{2}=(\mathcal{V}_{2},\mathcal{E}_{2})$ are isomorphic if there exists a bijection $f:\mathcal{V}_{1}\to\mathcal{V}_{2}$ such that a directed edge $(u,v)\in\mathcal{E}_{1}$ if and only if $(f(u),f(v))\in\mathcal{E}_{2}$. Furthermore, the repetition numbers of $(u,v)$ and $(f(u),f(v))$ remain the same if multiple edges or loops are allowed. This bijection $f$ is termed an isomorphism.
@@
In the reformulated problem, the graph isomorphism is used to exclude the symmetry of camera indices. In other words, swapping camera indices does not change the configuration. The edge-reversion symmetry is used to exclude the symmetry of two views. A directed graph $\mathcal{G}$ with edges =$\mathcal{E}=\{(u_{i},v_{i})\}_{i=1}^{n}$ and no isolated vertex can be represented by
@@
It is obvious that rearranging columns does not change the graph at all. In addition, a permutation of vertices in $E$ causes an isomorphic graph. Swapping two rows corresponds to reversing the edge directions of this graph. In this paper, we write $E_{1}\cong E_{2}$ if $E_{2}$ can be obtained by rearrangement of columns, swapping rows, and swapping vertices of $E_{1}$.
@@
It can be verified that $E_{3}$ can be converted to $E_{2}$ by swapping vertices $1$ and $3$; $E_{6}$ can be converted to $E_{2}$ by swapping vertices $1$ and $2$. $E_{5}$ can be converted to $E_{4}$ by swapping its two rows. As a result, there are $3$ distinct graphs. The distinct graphs with isomorphism and edge-reversion symmetry are $E_{1}$, $E_{2}\cong E_{3}\cong E_{6}$, and $E_{4}\cong E_{5}$.
0.D.2 A Combinatorics Solution to Enumeration
The problem is related to combinatorics and group theory. The Pólya enumeration theorem can be applied to solve this problem [17, 41]. To apply this theorem, we need to generate the cycle index of the permutation group $S_{n}$ acting on $n$ elements. Then we obtain the number of isomorphic graphs from this cycle index of $S_{n}$.
@@ 0.D.2 A Combinatorics Solution to Enumeration
0.D.3 Statistics of Cases and Solvers
Listing 2: Enumerate Distinct Graphs Using Matlab
@@ 0.D.3 Statistics of Cases and Solvers
Table 5: The number $D_{n}$ of distinct graphs given $n$ edges.
#edge $n$ 1 2 3 4 5 6 7 8 9 10
#distinct graph 2 9 37 186 985 5953 38689 271492 2016845 15767277
#camera 1 2 3 4 5 6 7 8 9 10 11 12
#cases 1 29 270 1029 1776 1630 853 280 66 15 3 1
#camera 1 2 3 4 5 6 7 8 9 10 11 12 13 14
#cases 1 39 568 3316 8599 11516 8787 4170 1296 312 66 15 3 1
@@
  • Match type $6\cup\phi$: All the 6 edges are same. There are $2$ cases belonging to this match type, including
    $E_{1}=\left[\begin{array}[]{cccccc}1&1&1&1&1&1\\ 1&1&1&1&1&1\end{array}\right],$
    $E_{2}=\left[\begin{array}[]{cccccc}1&1&1&1&1&1\\ 2&2&2&2&2&2\end{array}\right].$
    These two cases essentially correspond to monocular cameras. Since $6$ PCs provide an over-determined equation system, the equation system has no solution.
    $E_{3}=\left[\begin{array}[]{cccccc}1&1&1&1&1&2\\ 1&1&1&1&1&3\end{array}\right].$
    $E_{4}=\left[\begin{array}[]{cccccc}1&1&1&1&2&4\\ 1&1&1&1&3&5\end{array}\right].$
  • Match type $3\cup 3$: The maximal repetitive number of multiple edges is $3$, and there are two groups of multiple edges satisfying this condition. For example, the previously defined inter-camera and intra-camera cases belong to this match type. They are
    $E_{5}=\left[\begin{array}[]{cccccc}1&1&1&2&2&2\\ 2&2&2&1&1&1\end{array}\right],$
    $E_{6}=\left[\begin{array}[]{cccccc}1&1&1&2&2&2\\ 1&1&1&2&2&2\end{array}\right].$
  • Match type $3\cup x$: The maximal repetitive number of multiple edges is $3$, and there is only one group of multiple edges satisfying this condition. For example,
    $E_{7}=\left[\begin{array}[]{cccccc}1&1&1&2&4&6\\ 1&1&1&3&5&7\end{array}\right].$
  • Match type $\cup_{i}x_{i}$: The maximal repetitive number of multiple edges is less than or equal to $2$. For example,
    $E_{8}=\left[\begin{array}[]{cccccc}1&3&5&7&9&11\\ 2&4&6&8&10&12\end{array}\right].$
For different match types, the numbers of equations in $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ are different. We found that cases belonging to the same match type have preciously the same number of solutions. Tab. 8 shows the number of solutions for different match types in six-point problems for multi-camera systems.
Appendix 0.E Degenerate Configurations
For two-camera rigs, there are degenerate cases for 6pt+cayl/quat+generic. The devised solvers 6pt+cayl/quat+inter and 6pt+cayl/quat+intra have better performance than the generic solver in those cases. We also introduce two degenerate cases for these two solvers for two-camera rigs: (1) For inter-camera case, a two-camera rig undergoes pure translation while the baselines of the two cameras are parallel to the translation direction, (2) For intra-camera case, a two-camera rig undergoes pure translation or the cameras move along concentric circles. In these critical configurations, the metric scale of the translation is unobtainable.
Appendix 0.F Experiments
@@
In this section, we add more experiments about the 6DOF relative pose estimation of a multi-camera system. As described in the paper, all the solvers are implemented in C++. The source codes for 17pt-Li [36] and 8pt-Kneip [24] are adopted from the OpenGV library [23]. The source codes for 6pt-Stewénius [48] is adopted from the PoseLib library [30]. The efficiency and numerical stability of the proposed solvers are shown in Sec. 0.F.1. Expected runtime vs. inlier ratio accounting for numerical stability is demonstrated in Sec. 0.F.2. More experiments on synthetic data are shown in Sec. 0.F.3. The real-world experiments on KITTI and nuScenes datasets are shown in Sec. 0.F.4 and Sec. 0.F.5, respectively.
0.F.1 Efficiency and Numerical Stability
The runtimes of the solvers for multi-camera systems are assessed using an Intel(R) Core(TM) i7-7800X 3.50 GHz processor. The average runtime of the solvers over $10,000$ runs is shown in Tab. 9. As a linear solver, 17pt-Li is the most efficient one. All the proposed solvers have runtimes of $0.8\sim 1.8$ milliseconds. The solvers 6pt-Our-inter56, 6pt-Our-inter48, and 6pt-Our-intra are more efficient than the seminal 6pt-Stewénius among the minimal solvers.
Runtime 43.3 102.0 1214.9 1724.6 1053.4 827.8 774.1
(a) Revolution error $\varepsilon_{R}$.
(b) Translation error $\varepsilon_{\mathbf{t}}$.
(c) Translation dir. error $\varepsilon_{\mathbf{t},\text{dir}}$.
Figure 7: Probability density functions of the relative pose estimation error on noise-free observations under the first scenario. We designed a simulated two-camera rig comprising two perspective cameras. The orientations of two perspective cameras are roughly forward-facing with random perturbation. The horizontal axis corresponds to the $\log_{10}$ estimated errors, and the vertical axis is the empirical probability density.
(a) Rotation error $\varepsilon_{\mathbf{R}}$.
(b) Translation error $\varepsilon_{\mathbf{t}}$.
(c) Translation dir. error $\varepsilon_{\mathbf{t},\text{dir}}$.
(d) Rotation error $\varepsilon_{\mathbf{R}}$.
(e) Translation error $\varepsilon_{\mathbf{t}}$.
(f) Translation dir. error $\varepsilon_{\mathbf{t},\text{dir}}$.
(g) Rotation error $\varepsilon_{\mathbf{R}}$.
(h) Translation error $\varepsilon_{\mathbf{t}}$.
(i) Translation dir. error $\varepsilon_{\mathbf{t},\text{dir}}$.
(j) Rotation error $\varepsilon_{\mathbf{R}}$.
(k) Translation error $\varepsilon_{\mathbf{t}}$.
(l) Translation dir. error $\varepsilon_{\mathbf{t},\text{dir}}$.
Figure 8: Probability density functions of the relative pose estimation error on noise-free observations under the second scenario. We designed a simulated multi-camera system comprising twelve omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. The horizontal axis corresponds to the $\log_{10}$ estimated errors, and the vertical axis is the empirical probability density.
@@ 0.F.1 Efficiency and Numerical Stability
To test the numerical stability, we design two simulated scenarios. First, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras. Second, we design a simulated generalized camera composed of twelve omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random.
Fig. 7 illustrates the numerical stability of the solvers when applied to noise-free observations under the first scenario. The procedure is repeated for $10,000$ times on both inter-camera and intra-camera cases. The resulting empirical probability density functions are represented as a function of the $\log_{10}$ estimated errors. The numerical stability of a solver is determined by many factors, such as the problem’s complexity, parameterization, equation system construction, solver generator, and implementation. Usually, more efficient solvers generated by effective methods have less round-off error and thus have better numerical stability. Among these solvers, the 17pt-Li solver demonstrates the best numerical stability due to its status as a linear solver with fewer computations. Since the 8pt-Kneip solver relies on iterative optimization, it is susceptible to becoming trapped in local minima. The 6pt-Stewénius solver does not work well with a simulated two-camera rig. This phenomenon is consistent with [22], which observes the 6pt-Stewénius solver does not work for most axial cameras. All the proposed solvers exhibit satisfactory numerical stability. Since the 6pt-Our-inter56 solver displays better numerical stability than the 6pt-Our-inter48 solver, we recommend 6pt-Our-inter56 as the default solver for inter-camera case in the follow-up experiments.
@@
Fig. 8 illustrates the numerical stability of the solvers when applied to noise-free observations under the second scenario. The numerical stability of the solvers is tested in generic-camera, inter-camera, and intra-camera configurations. The strategy of correspondence selection is described in Sec. 0.F.3.2. For all three cases, the proposed solvers have satisfactory numerical stability. Among them, the 6pt-Our-inter56 solver has the best numerical stability. None of the comparison solvers can have satisfactory results for all the cases. Specifically,
@@
  • 6pt-Stewénius has satisfactory results for generic-camera configuration. Since this solver does not work for most axial cameras [22], it does not work well for inter-camera and intra-camera cases.
@@
Figure 9: The relationship between the RANSAC iteration number and the outlier ratio in the context of achieving a success probability of $99\%$.
0.F.2 Expected Runtime vs. Inlier Ratio Accounting for Numerical Stability
The computational time of RANSAC is contingent upon both the runtime of the minimal solver and the number of iterations. For the RANSAC methods, the iteration number is determined by the probability of choosing at least one inlier set. This calculation implicitly assumes that an accurate solution can be obtained given an inlier set. However, this is not always true when considering the numerical stability of the minimal solvers.
$(1-(1-\epsilon)^{s})^{N}=1-p,$ (19)
and
$N=\frac{\log(1-p)}{\log(1-(1-\epsilon)^{s})}.$ (20)
$(1-(p_{2}(1-\epsilon))^{s})^{\hat{N}}=1-p,$ (21)
and
$\hat{N}=\frac{\log(1-p)}{\log(1-(p_{2}(1-\epsilon))^{s})}.$ (22)
The left question is how to determine $p_{2}$ for a given minimal solver. For each minimal solver, we can empirically obtain the probability that the estimation errors $\varepsilon_{\mathbf{R}}$ and $\varepsilon_{\mathbf{t}}$ on noise-free observations are below specified thresholds. In this paper, we set the thresholds as $10^{-3}$. For the solvers 17pt-Li [36], 8pt-Kneip [24], 6pt-Stewénius [48], 6pt-Our-inter56, 6pt-Our-inter48 and 6pt-Our-intra, $p_{2}$ can be computed as 1.00, 0.09, 0.61, 0.99, 0.90 and 0.81, respectively. Fig. 9 shows RANSAC iteration numbers with respect to the outlier ratio for success probability $99\%$.
0.F.3 Experiments on Synthetic Data
Following the experiments about numerical stability, two simulated scenarios are designed and tested for the synthetic experiments, including scenarios for a two-camera rig and a generalized camera. The proposed solvers are evaluated against state-of-the-art solvers including 17pt-Li [36], 8pt-Kneip [24], and 6pt-Stewénius [48].